OpenStudy (cutiecomittee123):

HELP ME FILL IN THE BLANK FOR THIS EQUATION!! sin pi/4 sin pi/6 = 1/2 (cos pi/12 "blank" 5pi/12)

OpenStudy (cutiecomittee123):

btw I know we have to simplify! so sin pi/4 = sqrt2/2 sin pi/6 = 1/2 and cos pi/12 = a little harder to find lol

OpenStudy (ash2326):

We have \[\sin {\frac{\pi}{4}}\times \sin{\frac{\pi}{6}}=\frac{1}{2}(\cos{\frac{\pi}{12}}\times ..............(\frac{5\pi}{12})\]

OpenStudy (ash2326):

do you know the relation of half angle formula? for cos

OpenStudy (cutiecomittee123):

no x between the two sins on the left side!

OpenStudy (ash2326):

it's multiply sign

OpenStudy (cutiecomittee123):

Uhh I think its in my notes, oh okay:)

OpenStudy (ash2326):

go through your notes and try to find the formula. If you aren't able to, I will help..

OpenStudy (cutiecomittee123):

cos(x/2)=sqrt((1+cosx)/(2))

OpenStudy (cutiecomittee123):

right?

OpenStudy (ash2326):

good, now we can use this here to find cos (pi/12) we know cos (pi/6) so: \[cos {\frac{\pi}{12}}=\sqrt{\frac{1+cos {\frac{\pi}{6}}}{2}}\] could you substitute the values and find cos 15 degrees?

OpenStudy (cutiecomittee123):

cos pi/12 = sqrt ((1+sqrt3/2) / (2))

OpenStudy (ash2326):

let's simplify it a bit \[cos {\frac{\pi}{12}}=\sqrt {\frac{2+\sqrt3}{4}}\] ok, could you put this in the given expression and simplify?

OpenStudy (cutiecomittee123):

did you just multiply everything by 2? Hmm I am a little lost, so now the "cospi/12" will be replaced with the equation we just came up with?

OpenStudy (ash2326):

yep, I just multiplied both numerator and denominator in square root by 2. yes, we need to plugin the value in the question given

OpenStudy (cutiecomittee123):

\[ ( \sqrt2/2) (1/2) = \sqrt(2+\sqrt3)/4) blank 5\pi/12\]

OpenStudy (cutiecomittee123):

\[\sqrt2/4 = \sqrt(2+3)/(4) blank 5\pi/12\]

OpenStudy (cutiecomittee123):

oops that 3 should be sqrt3!

OpenStudy (ash2326):

correct so we have \[\frac{\sqrt 2}{4}=\frac{\sqrt{2+\sqrt 3}}{2} X\] I have replaced blank 5pi/12 by X

OpenStudy (cutiecomittee123):

alright, good idea!

OpenStudy (ash2326):

Oh i didn't notice, but let's go ahead and can you solve this further: have X one side and other things on other side

OpenStudy (cutiecomittee123):

shouldnt the number under the fraction on the right side be 4?

OpenStudy (cutiecomittee123):

so divide both sides by the coeeficient on the right side.

OpenStudy (ash2326):

I took the 4 out of square root, notice only the numerator is in the square root...

OpenStudy (cutiecomittee123):

oh i see!

OpenStudy (ash2326):

yes, divide both sides by the right side expression

OpenStudy (cutiecomittee123):

Yeah that is really difficult, I am stumped on how to divide the left side by such a weird number

OpenStudy (ash2326):

okay, let me show you. We will work in step by step manner \[\frac{\sqrt 2}{4} \times \frac{2}{\sqrt{2+\sqrt 3}}= X\] do you follow this?

zepdrix (zepdrix):

I would probably recommend using the Cosine Angle Addition/Subtraction Formula instead of the Half-Angle, so you don't end up with nested roots like that.\[\large\rm \cos\left(\frac{\pi}{12}\right)\quad=\cos\left(\frac{4\pi}{12}-\frac{3\pi}{12}\right)\quad=\cos\left(\frac{\pi}{3}-\frac{\pi}{4}\right)\]\[\large\rm =\cos\left(\frac{\pi}{3}\right)\cos\left(\frac{\pi}{4}\right)+\sin\left(\frac{\pi}{3}\right)\sin\left(\frac{\pi}{4}\right)\] I'm still really confused by this problem though. Is the "blank" referring to an operation? Or a trig function? Is the 5pi/12 supposed to be included in the cosine angle on the right?

zepdrix (zepdrix):

This problem doesn't make a lot of sense. Do you have a screen shot or something?

OpenStudy (ash2326):

@zepdrix blank is supposed to be a trigonometric function..

OpenStudy (cutiecomittee123):

Yes @ash2326 I think that is correct, the blank should be a trig function.

OpenStudy (ash2326):

@cutiecomittee123 Did you follow my last step?

zepdrix (zepdrix):

Oh, there is a minus in front of the blank...

OpenStudy (cutiecomittee123):

which doesnt necessarily change anything thus far? right?

hartnn (hartnn):

i think they were expecting us to use \(\Large \cos y - \cos x = 2 [\sin \dfrac{(x+y)}{2} \sin \dfrac{(x-y)}{2}] \)

hartnn (hartnn):

I didn't calculate, but i think with (x+y)/2 = pi/4 and (x-y)/2 = pi/6 we would get x = pi/12 and y = 5pi/12

hartnn (hartnn):

oops, we actually get x = 5pi/12 and y = pi/12

hartnn (hartnn):

got what i did there?

OpenStudy (ash2326):

@hartnn is right, I will let him help you @cutiecomittee123 I will be here to help in case it's needed but I don't think that would be the case

OpenStudy (cutiecomittee123):

okay I am trying to follow, it is a little challenging but I think I got it so far.

OpenStudy (cutiecomittee123):

I dont know what the equation that you represented for us to use is?

hartnn (hartnn):

since we had this formula \( \cos y - \cos x = 2 [\sin \dfrac{(x+y)}{2} \sin \dfrac{(x-y)}{2}]\) and we had a 2[sin pi/4 sin pi/6 ] so comparing we can say that (x+y)/2 = pi/4 (x-y)/2 = pi/6 once we get the value of x and y we can substitute in cos y - cos x to get our final answer

hartnn (hartnn):

solving (x+y)/2 = pi/4 (x-y)/2 = pi/6 is purely algebraic

OpenStudy (cutiecomittee123):

but how is this relevant to our origional equation?

hartnn (hartnn):

the original question has similar form as that of the formula

hartnn (hartnn):